OpenStudy (anonymous):

Which number is halfway between 3/5 and 7/10? A.27/40 B.2/3 C.7/10 D.13/20 idk plz help! Medal

OpenStudy (anonymous):

@ayyookyndall @mathmate @Vocaloid

OpenStudy (anonymous):

This is the last one

OpenStudy (anonymous):

their all between .6 and .7 which are the decimals of 3/5 and 7/10

OpenStudy (anonymous):

@ayyookyndall

OpenStudy (mathmate):

Example: The number between \(\large \frac{4}{9}\) and \(\frac{2}{7}\) can be found by adding them and dividing by two. So \(\huge \frac{\frac{4}{9} +\frac{2}{7}}{2}=\frac{\frac{28}{63} +\frac{12}{63}}{2}=\frac{\frac{40}{63}}{2}=\frac{20}{63}\)

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

i still dont fully get it srry

OpenStudy (anonymous):

how did the first to become demonaters of 63?

OpenStudy (anonymous):

wait did you multiply? @mathmate

OpenStudy (anonymous):

ok i get that now lol but how did you get the numeraters?

OpenStudy (anonymous):

you there?

OpenStudy (mathmate):

Yes, but having difficulties with the connection!

OpenStudy (anonymous):

oh srry

OpenStudy (mathmate):

28/63 is an equivalent fraction of 4/9, but with a denominator of 63.